# Benchmarking Android Malware Detection: Rethinking the Role of Traditional and Deep Learning Models

### v1

@@ Abstract
Android malware detection has been extensively studied using both traditional machine learning (ML) and deep learning (DL) approaches. While many state-of-the-art detection models, particularly those based on DL, claim superior performance, they often rely on limited comparisons, lacking comprehensive benchmarking against traditional ML models across diverse datasets. This raises concerns about the robustness of DL-based approaches’ performance and the potential oversight of simpler, more efficient ML models. In this paper, we conduct a systematic evaluation of Android malware detection models across four datasets: three recently published, publicly available datasets and a large-scale dataset we systematically collected. We implement a range of traditional ML models, including Random Forests (RF) and CatBoost, alongside advanced DL models such as Capsule Graph Neural Networks (CapsGNN), BERT-based models, and ExcelFormer-based models. Our results reveal that while advanced DL models can achieve strong performance, they are often compared against an insufficient number of traditional ML baselines. In many cases, simpler and more computationally efficient ML models achieve comparable or even superior performance. These findings highlight the need for rigorous benchmarking in Android malware detection research. We encourage future studies to conduct more comprehensive benchmarking comparisons between traditional and advanced models to ensure a more accurate assessment of detection capabilities. To facilitate further research, we provide access to our dataset, including app IDs, hash values, and labels.
@@ 1 Introduction
The Android ecosystem maintains the leading position as the largest and most active system among the worldwide mobile computing environment by far. Among its ecosystem components, the app markets which consist of Google Play, as well as manufacturer and third-party app stores, play a crucial role in releasing and distributing the apps to end users. According to one recent 42matters’s statistics [1], there were 40,325 apps on average released monthly just through Google Play Store between November 2024 and Juanuary 2025. The app market’s rapid growth makes it crucial to have an automated and effective app vetting system that scales to the app market sizes. An important component of a vetting system is the application of machine learning to triage the apps in the vetting process.
Traditional machine learning approaches (shorthanded ML hereafter) such as Naive Bayes (NB), k-Nearest Neighbors (k-NN), Support Vector Machines (SVM), Random Forest (RF), Artificial Neural Networks (ANN), Gradient Boosting, etc. have been applied to train models on a training dataset, which consists of apps whose maliciousness is known. Then the trained models are applied to classify a test dataset, where apps’ maliciousness is unknown. Many traditional ML models have shown outstanding performance [2, 3, 4, 5, 6, 7, 8, 9, 10, 11] on datasets from small to market-scale.
@@
More recently, deep learning approaches (shorthanded DL hereafter) have been studied on Android malware detection [12, 13, 14, 15, 16, 17, 18, 19, 20, 21, 22, 23, 24, 25, 26]. One potential advantage of DL over ML is that it does not require an extensive feature selection and engineering process, although careful consideration still needs to be applied as to how to represent app data provided as input to the DL models.
Despite the proposed DL approaches for this problem, there exist limited comparisons [12, 21, 27, 24, 26] between ML and the proposed DL models. More often than not researchers do not publish the dataset used to compare performances of the various models, making it impossible for other researchers to further benchmark the model against additional alternatives. It is encouraging that we are seeing some recent works where researchers publish the datasets used in their study. Upon looking through the literature, we identified two works where 1) the researchers compared novel DL models against ML models when applied to Android malware detection, 2) the researchers published the datasets used in their experiments, and 3) the datasets are of substantial sizes. This enabled us to conduct additional experiments to benchmark the proposed DL models against additional ML models that were not examined in the original research. In addition, we conducted benchmarking of two other DL models against the ML models, on a large dataset that we curated. Our experiments show mixed results regarding whether DL models can outperform ML models for Android malware detection. The results call for more extensive benchmarking of new deep learning models against traditional machine learning models for Android malware detection.
2 Related Work
Machine learning-based Android malware detection systems have been researched for a decade. Crowdroid [2] is an earlier ML framework for Android malware detection. DroidAPIMiner [3] performs static analysis to extract lightweight features at API level, then uses different classification models such as k-NN, SVM and Decision Trees (DT) to detect malware. DREBIN [6] uses similar static analysis techniques as DroidAPIMiner to generate a feature set that contains components, permissions, filtered intents, API calls and network addresses, then uses SVM as a malware classifier. Roy et al. [28] performed experimentation studies to investigate a number of research questions related to using ML in Android malware detection. MaMaDroid [29] abstracts apps’ API calls from 340 packages and 11 families of malware, then applies a Markov chain to form the sequence of abstracted API calls extracted from call graphs as a feature vector, and finally performs classification using Random Forest, k-NN and SVM algorithms. APIChecker [11] uses dynamic analysis to collect API features, applies nine machine learning algorithms for Android malware detection and evaluates their performance on the T-Market data with about 500K apps. The results show that RF yields the best overall performance among the selected models. DroidAPIMiner [4], DREBIN [6], DroidMat [30], and MaMaDroid [29] also compared their performance with other traditional ML classifiers and anti-virus scanners. MaMaDroid uses four different ML algorithms and the detection results obtained with RF have better performance in general.
Recent years have seen an increase in deep learning related Android malware detection research. The deep neural network models used include deep belief networks (DBN) [12, 13], convolutional neural networks (CNN) [15, 16, 24, 31], pre-trained BERT models [26], recurrent neural networks (mainly long short-term memory networks, or LSTM) [32, 19], and multilayer perceptrons (MLP) with continuous learning schemes [25], among others. Droid-Sec [12] trained and tested a DBN model based on 500 apps (250 malware). Experimental results showed that the DL model has significant accuracy improvement as compared to traditional ML models, such as SVM, NB, Logistic Regression, and MLP. Deep4MalDroid [14] evaluated a Stacked AutoEncoders (SAEs) model based on 3000 apps (50% of which are malware). Results showed that the proposed DL model has higher accuracy than the selected ML models, specifically, SVM, ANN, NB, and DT. DeepFlow [33] proposed a DBN model, and conducted experiments based on 3000 benign apps and 8000 malicious apps. The comparison between DL and traditional ML models showed that DBN and SVM outperform other selected ML models in balancing the recall and precision. DexRay [24] converts the raw bytecode of DEX files in the APK as one stream vector, then converts the vector to a grey-scale image which is provided as input to a CNN architecture. DexRay evaluated the performance of the CNN model based on 61,809 malicious and 134,134 benign apps, then compared its detection results with DREBIN’s results and other image-based malware detection approaches. DexRay yields almost the same performance metrics as DREBIN; however, its performance can be significantly affected if obfuscated apps are absent in the training data but present in the test data. DetectBERT [26] leverages a pre-trained BERT-like model with correlated Multiple Instance Learning (c-MIL) to process Smali code from APKs, enriching class-level features and aggregating them into an app-level representation for effective Android malware detection. It employs the DexRay dataset as its benchmark baseline for evaluation. DeepRefiner [17] is a two-layer detection system, with the first detection layer based on MLP, and the second layer based on an LSTM network. The experimental data include 47,525 benign apps and 62,915 malicious apps. The detection performance, based on accuracy, TPR and FPR, is approximately 10% higher than that of a fine-turned SVM model. DANdroid [34] uses the DREBIN dataset to evaluate its proposed DL discriminative adversarial network, and the performance is very close to that of DREBIN. DL-Droid [23] generates 420 dynamic stateful features such as application attributes, actions and permissions by running apps on real devices, then feeding them into a multilayer perceptron classifier. It uses weighted F measure (w-FM) as main evaluation metric and compares it against that of seven selected ML algorithms. DL-Droid reached the highest w-FM value of 98.82% (the best ML model RF reached w-FM value of 97.1%). AMD-CNN [31] is a visualization-based Android malware detection tool. It extracts the features from Android manifest file and converts them to RGC images, then feeds the images to a CNN network. It was evaluated based on 960 benign and 960 malicious apps. [35] identified five factors that influence the training and performance of ML-based Android malware detectors. The authors conducted a fair comparison of eight detectors on a mixed-market dataset, accounting for these factors. The detectors include seven traditional ML models (across four types) and one DBN with two hidden layers of 150 neurons each. Chen et al. [25] proposed a continuous learning approach using MLP to enhance the adaptability of ML-based Android malware detectors to concept drift [36]. Their work includes a comprehensive comparison with major baseline models, with both input data and model-related code available as open source. The evaluation was conducted using two established datasets, APIGRAPH [37] and AndroZoo [38].
Our study complements these prior works by extensive benchmarking the proposed DL models against additional standard ML models that were not examined by the original researchers. Our results lead to reassessment of DL’s advantage against ML in Android malware detection, and highlights the need for extensive benchmarking for this problem domain.
3 Benchmarking Methodology
@@
Benchmarking across different models is a standard approach for evaluating and comparing their performance, as well as identifying the strengths and weaknesses of various approaches. However, determining the sufficient number of models to compare with is not easy, as further benchmarking using more models could always be beneficial. For this reason, it is utterly important that researchers make the dataset upon which their models are evaluated available. This will enable other researchers to evaluate additional models that had not been considered in the original researcher. More extensive evaluations can provide deeper insights and help understand the true strengths of the various models for Android malware detection.
One of the key challenges for such extensive evaluation is that not all works publish the datasets used in their study. Fortunately, we have identified recent studies, such as Chen et al.’s Enc+MLP [25] and Sun et al.’s DetectBERT [26], that have made their datasets publicly available. This allowed us conduct more extensive benchmarking of the models proposed in these two works. Additionally, we collected a Google Play-only dataset and used two DL approaches, Capsule Graph Neural Networks (CapsGNN) and ExcelFormer, which have not been explored before for Android malware detection. Each of the three works employs different datasets to evaluate their models, and their approaches to app representation also vary. Our primary objective is not to assert the superiority of any specific model but to demonstrate how comprehensive benchmarking enhances the understanding of model performance. To this end we adopt the following benchmarking strategy.
@@
In the following sections, we first discuss traditional machine learning models considered and then explore additional models that were not included in the two prior studies.
4 Traditional Machine Learning Models Considered
Widely explored ML models in Android malware detection include Naive Bayes (NB) [39, 40], support vector machines (SVM) [4, 6], k-nearest neighbors (k-NN) [30, 40], and random forest (RF) [29]. We consider these four widely used models as potential baselines to compare with the proposed DL models. Additionally, Gradient Boosted Decision Trees (GBDTs) are powerful classification tools, with CatBoost [41] being a notable variant of GBDT-based ensemble techniques. Having been introduced in late 2018, CatBoost has demonstrated strong performance in classification and regression tasks across various domains, including cybersecurity [41]. Thus we also include it as one of the ML models in our benchmarking.
NB: This is a simple but effective classifier. Furthermore, it is the fastest classifier we evaluated. It can get the classification result quickly even with half a million apps.
SVM: Widely used classifier in Android malware classification. It requires more memory and computation time during the training stage, which poses a significant challenge for large datasets. We set the kernel to be linear, and use a grid search with values 0.001, 0.1, 1.0, and 10 to decide the optimal regularization parameter to train the SVM model.
k-NN: Another simple and robust traditional classifier. The value of the hyper-parameter $k$ has a direct impact on the performance. We use a grid search with values between 3 and 15 to determine the optimal $k$ value to train the k-NN model.
RF: A popular ensemble-type model based on decision trees. It has shown promising performance in Android malware detection [29] and has the capability to be applied to market-scale datasets [11]. We perform a grid search to find the optimal number of trees in the forest. The estimator value is between 100 and 300.
CatBoost: Another ensemble-based model built on decision trees. While it is relatively new compared to the other ML models, recent studies across various domains have demonstrated its effectiveness in classification tasks [41]. CatBoost is sensitive to hyperparameters, and our experimental results are based on the following setup: 1000 iterations, a learning rate of 0.1, a depth of 10, and an L2 regularization parameter (l2_leaf_reg) of 5.
5 Additional Benchmarking for DetectBERT
@@ 5.3 Additional Benchmarking
TABLE I: Performance comparison of various models on DetectBERT benchmark dataset. The unshaded are from the original paper, while the shaded results are from our benchmarking.
@@ 6.2 Datasets for Enc+MLP
APIGraph Dataset. Collected by Chen et al. [25] using the app hash list provided by APIGraph [37], this dataset spans seven years of Android apps from 2012 to 2018. The apps are ordered based on their appearance timestamps in VirusTotal [42], addressing both spatial and temporal biases [43]. The dataset is evenly distributed across the years, with each year consisting of approximately 90% benign apps and 10% malicious apps. The malicious apps are sourced from VirusTotal [42], VirusShare [44], and the AMD dataset [45], while the benign apps are obtained from AndroZoo [38]. Table III provides a detailed breakdown of this dataset.
6.3 Additional Benchmarking
Monthly Sample Budget Model Architecture Sample Selector APIGraph Dataset Average Performance (%) AndroZoo Dataset Average Performance (%)
@@ 6.3 Additional Benchmarking
We selected additional ML models from the ones mentioned in Section 4: NB, KNN, RF, and CatBoost (excluding SVM, as it was already included as a baseline in the original work), to expand the set of baseline models. Table IV presents the performance of all tested models, with shaded rows indicating models not included in the original work. If an additional model did not outperform the original baseline models, we omitted its results from the table. As a result, Table IV only includes RF and CatBoost as additional baseline models. The results show that RF achieves performance comparable to the ‘Enc+MLP’ model on the APIGraph dataset but performs poorly on the AndroZoo dataset. However, CatBoost outperforms ‘Enc+MLP’ on both datasets across different labeling budgets, except for the AndroZoo dataset when the labeling budget is set to 400.
7 Benchmarking for CapsGNN
@@
7.1 Benchmarking for CapsGNN
CapsGNN is a DL model designed to capture complex patterns within an app. We include five widely used ML models, as mentioned in Section 4, as potential baselines for comparison with the CapsGNN model.
@@ 7.1 Benchmarking for CapsGNN
Traditional ML models usually need a predefined feature set to represent an application, and the performance of the classifier highly depends on the features. Section 2 briefly explains some prior works’ approaches to feature selection. The most common features used in prior works are API calls, app permissions, and a few other pieces of information from raw app code. The feature sources are similar but different approaches collect and organize the features in different ways. Among prior works based on traidtional ML, DREBIN [6] has shown high detection preformance on datasets with realistic size and malware to benign app ratio. Recent work [51] shows that DREBIN’s features are still robust for apps as late as 2019. DREBIN extracts static features such as components, intents, API calls, permissions, and network addresses totaling 500,000 features. Roy et al. [28] achieved similar performance as DREBIN by choosing only 471 static features from the original DREBIN’s feature set. To balance the computational cost and performance on large-scale datasets, we use the same features as those in Roy et al.’s work [28], which consist of 471 features including permissions, intent actions, discriminative APIs, obfuscation signatures, and native code signatures. To support future research, the complete feature list will release as open-source with other related code following the acceptance of this work.
7.2 Dataset for CapsGNN
@@
Figure 3: Top-performing models among all tested models

### v2

# Benchmarking Android Malware Detection: Traditional vs. Deep Learning Models

@@ Abstract
Android malware detection has been extensively studied using both traditional machine learning (ML) and deep learning (DL) approaches. While many state-of-the-art detection models, particularly those based on DL, claim superior performance, they often rely on limited comparisons, lacking comprehensive benchmarking against traditional ML models across diverse datasets. This raises concerns about the robustness of DL-based approaches’ performance and the potential oversight of simpler, more efficient ML models. In this paper, we conduct a systematic evaluation of Android malware detection models across four datasets: three recently published, publicly available datasets and a large-scale dataset we systematically collected. We implement a range of traditional ML models, including Random Forests (RF) and CatBoost, alongside advanced DL models such as Capsule Graph Neural Networks (CapsGNN), BERT-based models, and ExcelFormer-based models. Our results reveal that in many cases simpler and more computationally efficient ML models achieve comparable or even superior performance compared with DL models. These findings highlight the need for rigorous benchmarking in Android malware detection research. We encourage future studies to conduct more comprehensive benchmarking comparisons between traditional and advanced models to ensure a more accurate assessment of detection capabilities. To facilitate further research, we provide access to our dataset, including app IDs, hash values, and labels.
@@ I Introduction
The Android ecosystem maintains the leading position as the largest and most active system among the worldwide mobile computing environment by far. According to one recent 42matters’s statistics [1], there were 40,325 apps on average released monthly through Google Play alone between November 2024 and Juanuary 2025. The app market’s rapid growth makes it crucial to have an automated and effective app vetting system that scales to the app market sizes. An important component of a vetting system is the application of machine learning to triage the apps in the vetting process. Researchers train a machine learning model on a dataset consistsing of apps whose maliciousness is known. Then the trained model is applied to a disjoint test dataset to evaluate the model’s effectiveness.
@@
Traditional machine learning approaches (shorthanded ML hereafter) such as Naive Bayes (NB), k-Nearest Neighbors (k-NN), Support Vector Machines (SVM), Random Forest (RF), Artificial Neural Networks (ANN), Gradient Boosting, and so on, have shown outstanding performance [2, 3, 4, 5, 6, 7, 8, 9, 10, 11] on datasets from small to market-scale. More recently, deep learning approaches (shorthanded DL hereafter) have been studied on Android malware detection [12, 13, 14, 15, 16, 17, 18, 19, 20, 21, 22, 23, 24, 25, 26]. One potential advantage of DL over ML is that it does not require an extensive feature selection and engineering process, although careful consideration still needs to be applied as to how to represent app data provided as input to the DL models.
Despite various DL approaches proposed for Android malware detection, there is limited comparison between DL and traditional ML models [12, 21, 27, 24, 26]. Evaluating DL models with only a few baselines often fails to reflect true performance – a pitfall termed “inappropriate baseline” [28] previously noted in areas like vulnerability discovery and network intrusion detection. We observed similar shortcomings in Android malware detection research. Many studies do not release their datasets, making it difficult to benchmark against a broader range of alternatives. Encouragingly, some recent works have made their datasets publicly available. Through our literature review, we identified two studies that 1) compared DL and ML models, 2) published their datasets, and 3) used datasets of substantial sizes. This allowed us to further benchmark those DL models against additional ML models. We also evaluated two other DL models against ML models using a large dataset we curated. Our experiments show mixed results regarding whether DL consistently outperforms ML models, underscoring the need for more comprehensive benchmarking in Android malware detection.
II Related Work
Machine learning-based Android malware detection has been studied for over a decade. Early approaches like Crowdroid [2] used k-means clustering to separate benign apps from malware, DroidAPIMiner [3] extracted static API-level features and applied classifiers such as k-NN, SVM, and Decision Trees (DT). DREBIN [6] represented apps using a rich static feature set and classified them using SVM. MaMaDroid [29] abstracted API calls into Markov chains and used classifiers like RF and SVM. APIChecker [11] collected dynamic API features and evaluated nine ML algorithms, with RF performing best on a dataset of approximately 500K apps. Other systems, including DroidMat [30], APIGraph [31], and MalScan [32], employed lightweight static or graph-based features with classifiers like k-NN and RF, consistently demonstrating strong detection performance.
Recent research has increasingly focused on deep learning (DL) for Android malware detection, leveraging models such as deep belief networks (DBN) [12, 13], convolutional neural networks (CNN) [15, 16, 24, 33], long short-term memory (LSTM) networks [34, 19], pre-trained BERT models [26], and multilayer perceptrons (MLP) [25]. Droid-Sec [12] used DBN and showed improved accuracy over traditional ML models. Deep4MalDroid [14] applied stacked autoencoders (SAEs) and outperformed classifiers like SVM and NB. DeepFlow [35] showed DBN and SVM achieved better balance between recall and precision than other ML models. DexRay [24] transformed DEX bytecode into grayscale images for CNN-based detection, achieving performance comparable to DREBIN. DetectBERT [26] used BERT [36] with Multiple Instance Learning to process Smali code, improving class-level representation. DeepRefiner [17] combined MLP and LSTM in a two-layer system, achieving  10% higher accuracy than SVM. DANdroid [37] evaluated a discriminative adversarial network on the DREBIN dataset, showing similar results. DL-Droid [23] extracted 420 dynamic features and used MLP, outperforming seven ML baselines. AMD-CNN [33] visualized Android manifest features as RGC images for CNN input. Chen et al. [25] proposed a continuous learning MLP model for handling concept drift [38], evaluated on APIGraph [31] and AndroZoo [39] with major baseline models.
III Benchmarking Methodology
@@
Benchmarking across different models is a standard approach for evaluating and comparing their performance, as well as identifying the strengths and weaknesses of various approaches. However, determining the sufficient number of models to compare with is not easy, as further benchmarking using more models could always be beneficial. For this reason, it is utterly important that researchers make the dataset upon which their models are evaluated available. This will enable other researchers to evaluate additional models that had not been considered in the original research. More extensive evaluations can provide deeper insights and help understand the true strengths of the various models for Android malware detection.
Our primary objective is not to assert the superiority of any specific model but to demonstrate how comprehensive benchmarking enhances the understanding of model performance. To this end we adopt the following benchmarking strategy.
@@
III-A Traditional Machine Learning Models Considered
Widely explored ML models in Android malware detection include Naive Bayes (NB) [40, 41], support vector machines (SVM) [4, 6], k-nearest neighbors (k-NN) [30, 41, 32], and random forest (RF) [29, 32]. We consider these four widely used models as potential baselines to compare with the proposed DL models. Additionally, Gradient Boosted Decision Trees (GBDTs) are powerful classification tools, with CatBoost [42] being a notable variant of GBDT-based ensemble techniques. Having been introduced in late 2018, CatBoost has demonstrated strong performance in classification and regression tasks across various domains, including cybersecurity [42]. Thus we also include it as one of the ML models in our benchmarking.
III-B Deep Learning Models Considered
One of the key challenges in applying recent DL methods to Android malware detection is the lack of publicly available datasets in their studies. Fortunately, we have identified recent studies, such as Chen et al.’s Enc+MLP [25] and Sun et al.’s DetectBERT [26], that have made their datasets publicly available. This allowed us to conduct more extensive benchmarking of the models proposed in these two works. Additionally, we collected a Google Play-only dataset and used two DL approaches, Capsule Graph Neural Networks (CapsGNN) and ExcelFormer, which have not been explored before for Android malware detection. Each of the three works employs different datasets to evaluate their models, and their approaches to app representation also vary.
IV Case Studies
IV-A Case Study 1: DetectBERT
@@ IV-A3 Additional Benchmarking
TABLE I: Model performance on the DetectBERT benchmark dataset. Unshaded results are from the original paper; shaded results are from our benchmarking.
@@ IV-B2 Datasets for Enc+MLP
APIGraph Dataset. Collected by Chen et al. [25] using the app hash list provided by APIGraph [31], this dataset spans seven years of Android apps from 2012 to 2018. The apps are ordered based on their appearance timestamps in VirusTotal [43], addressing both spatial and temporal biases [44]. The dataset is evenly distributed across the years, with each year consisting of approximately 90% benign apps and 10% malicious apps. The malicious apps are sourced from VirusTotal [43], VirusShare [45], and the AMD dataset [46], while the benign apps are obtained from AndroZoo [39]. Table II provides a detailed breakdown of this dataset.
Monthly
Sample
Budget
Model
Architecture
Sample
Selector
APIGraph Dataset
Average Performance (%)
AndroZoo Dataset
Average Performance (%)
@@ IV-B3 Additional Benchmarking
We selected additional ML models from the ones mentioned in Section III-A: NB, KNN, RF, and CatBoost (excluding SVM, as it was already included as a baseline in the original work), to expand the set of baseline models. Table IV presents the performance of all tested models, with shaded rows indicating models not included in the original work. If an additional model did not outperform the original baseline models, we omitted its results from the table. As a result, Table IV only includes RF and CatBoost as additional baseline models. The results show that RF achieves performance comparable to the Enc+MLP model on the APIGraph dataset but performs poorly on the AndroZoo dataset. However, CatBoost outperforms Enc+MLP on both datasets across different labeling budgets, except for the AndroZoo dataset when the labeling budget is set to 400.
IV-C Case Study 3: CapsGNN
@@
IV-C1 Benchmarking for CapsGNN
CapsGNN is a DL model designed to capture complex patterns within an app. We include five widely used ML models, as mentioned in Section III-A, as potential baselines for comparison with the CapsGNN model.
@@ IV-C1 Benchmarking for CapsGNN
Traditional ML models usually need a predefined feature set to represent an application, and the performance of the classifier highly depends on the features. Section II briefly explains some prior works’ approaches to feature selection. The most common features used in prior works are API calls, app permissions, and a few other pieces of information from raw app code. The feature sources are similar but different approaches collect and organize the features in different ways. Among prior works based on traidtional ML, DREBIN [6] has shown high detection preformance on datasets with realistic size and malware to benign app ratio. Recent work [52] shows that DREBIN’s features are still robust for apps as late as 2019. DREBIN extracts static features such as components, intents, API calls, permissions, and network addresses totaling 500,000 features. Roy et al. [53] achieved similar performance as DREBIN by choosing only 471 static features from the original DREBIN’s feature set. To balance the computational cost and performance on large-scale datasets, we use the same features as those in Roy et al.’s work [53], which consist of 471 features including permissions, intent actions, discriminative APIs, obfuscation signatures, and native code signatures. To support future research, the complete feature list will release as open-source with other related code following the acceptance of this work.
IV-C2 Dataset for CapsGNN
@@
Figure 3: Top-performing models among all tested models
Figure 4: RF and CapsGNN performance over the sliding windows